A slate falls from the roof of a height-rise building. find how far it falls
a) in the first second b) in the first two seconds c) during the third second
1
Expert's answer
2017-11-08T13:45:06-0500
Slate falls under action of gravitational attraction of the Earth, so
ma = mg from the second law of Newton, then
a = g slate moves with constant acceleration, with zero initial velocity,
and we can find distance from S = gt^2/2, where g = 9.8 m/s^2
t = 1s, S = 9.8*1/2 = 4.9 m;
t = 2s, S = 19.6m;
t = 3s, S = 44.1m - 19.6m = 24.5m.